OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 6 x – y = 3

OpenStudy (anonymous):

by multiplying equation 2 by 2 2x-2y=6 then add the first equation with the second 6x=12 then x=2

OpenStudy (anonymous):

sustitude x in equation(1) 4(2)+2y=6 y=-1

OpenStudy (anonymous):

checking 4(2)+2(-1)=6 u can use any equation to check

OpenStudy (anonymous):

Ok. You need to learn how to do these problems. There are two ways: elimination and substitution. I'll only demonstrate one method because it's kind of a pain in the a**. With elimination, you look at the two equations and determine whether you can eliminate one of the terms by switching its sign; e.g: 4x-3y=8 4x-8y=3 In this problem, you can switch one of the "4x's" to "-4x". It then becomes: 4x-3y=8 -4x-8y=3 Now you can eliminate both of the 4x's from the equations, leaving you with the more simplified system of: -3y=8 -8y=3 Simplify one of the equations and you have the value of y, then substitute the y-value in one of the equations of the original system, simplify it and you'll have x. This method is called ELIMINATION.